# Tensor Product Attention Is All You Need

### v4

@@ C.8 Inference Time Decoding of Different Attention Mechanisms
TPA Total: $\quad 1568M+2080M=3648M$
For MQA and GQA, assuming $H=32$ query heads and $d_{h}=64$:
MQA/GQA Total: $\quad 2MHd_{h}=2\cdot M\cdot 32\cdot 64=4096M$
For MLA, with $H=32,d_{c}=256,d_{h}^{R}=32$:
MLA Logits: $\quad MH(d_{c}+d_{h}^{R})=M\cdot 32\cdot(256+32)=M\cdot 32\cdot 288=9216M$
MLA Value Agg.: $\quad MHd_{c}=M\cdot 32\cdot 256=8192M$
MLA Total: $\quad M\cdot 32\cdot(2\cdot 256+32)=M\cdot 32\cdot 544=17408M$
In this configuration, TPA requires fewer FLOPs ($3648M$) than MHA ($4096M$). The FLOPs for scaled logits in TPA ($1568M$) are less than MHA’s ($2048M$). The value aggregation FLOPs for TPA ($2080M$) are comparable to MHA’s ($2048M$).
This reduction in FLOPs, particularly when $R_{K}$ and $R_{V}$ are small (e.g., 1), combined with optimized memory access patterns in algorithms like FlashTPA Decoding, contributes to TPA’s competitive decoding speed, especially for long sequences. The actual end-to-end wall-clock speedup also depends on kernel fusion, hardware specifics, and the efficiency of einsum implementations, but the factorized formulation offers a clear path to reduced computational load.
Appendix D More on FlashTPA Decoding Algorithm
This section provides further details on the FlashTPA Decoding algorithm (see Algorithm 1), including the setup of factorized components.
Factorized Component Setup for FlashTPA Decoding. Let $B$ be the batch size, $N$ the number of query tokens (which is 1 for decoding, $N=1$), $M$ the current length of the KV cache, $H$ the number of attention heads, and $R_{Q}$ the rank of the query factorization. The feature dimensions for the factorized components are $D$ (for query $\mathbf{B}_{Q}$ and key $\bm{b}^{K}$) and $E$ (for value $\bm{b}^{V}$). Typically, $D$ and $E$ correspond to the head dimension $d_{h}$.
The query for the current token $\mathbf{x}_{t}$ (where $t$ is the current time step, $N=1$) is factorized as $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$, where $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{H\times R_{Q}}$ (derived from the input tensor ‘Aq’ of shape $(B,1,H,R_{Q})$ for the current token) and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times D}$ (similarly from ‘Bq’ of shape $(B,1,R_{Q},D)$).
The cached keys $\mathbf{K}_{s}$ and values $\mathbf{V}_{s}$ for past tokens $s\in[1,M]$ are stored in their factorized form. For clarity in the data flow diagram (Figure 2 in the main paper) and the pseudo-code in Algorithm 1, we present the case where cached keys and values use ranks $R_{K}=1$ and $R_{V}=1$. The general TPA formulation allows for $R_{K}>1$ and $R_{V}>1$ for cached items (by storing, for example, $\mathbf{A}_{K}(\mathbf{x}_{s})\in\mathbb{R}^{R_{K}\times H}$ and $\mathbf{B}_{K}(\mathbf{x}_{s})\in\mathbb{R}^{R_{K}\times D}$). The einsum operations detailed in Algorithm 1 can be extended to accommodate higher ranks for keys and values by including summations over these additional rank dimensions. For $R_{K}=1$ and $R_{V}=1$, the cached $s$-th key and value are: $\mathbf{K}_{s}=\mathbf{a}^{K}(\mathbf{x}_{s})\otimes\bm{b}^{K}(\mathbf{x}_{s})\quad\in\mathbb{R}^{H\times D}$, $\mathbf{V}_{s}=\mathbf{a}^{V}(\mathbf{x}_{s})\otimes\bm{b}^{V}(\mathbf{x}_{s})\quad\in\mathbb{R}^{H\times E}$. Here, $\mathbf{a}^{K}(\mathbf{x}_{s})\in\mathbb{R}^{H}$ (from the cached tensor $\mathbf{a}^{K}_{\text{cache}}$ of shape $(B,M,H)$), $\bm{b}^{K}(\mathbf{x}_{s})\in\mathbb{R}^{D}$ (from $\bm{b}^{K}_{\text{cache}}$ of shape $(B,M,D)$), $\mathbf{a}^{V}(\mathbf{x}_{s})\in\mathbb{R}^{H}$ (from $\mathbf{a}^{V}_{\text{cache}}$ of shape $(B,M,H)$), and $\bm{b}^{V}(\mathbf{x}_{s})\in\mathbb{R}^{E}$ (from $\bm{b}^{V}_{\text{cache}}$ of shape $(B,M,E)$).
D.1 Detailed Computation Steps of FlashTPA Decoding Algorithm
The algorithm proceeds through a series of einsum operations, which correspond to specific tensor contractions:
  1. 1.
    Compute $\mathbf{B}_{Q}$-$\bm{b}^{K}$ inner products ($S^{(1)}$): For each query (batch $b$, query index $n=0$), each query rank component $r$, and each cached key position $m$, we compute the dot product of the $D$-dimensional feature vector of the $r$-th query factor $(\mathbf{B}_{Q})_{b,n,r,:}$ with the $D$-dimensional feature vector of the $m$-th cached key $(\bm{b}^{K}_{\text{cache}})_{b,m,:}$.
    $S^{(1)}_{b,n,m,r}=\sum_{d=1}^{D}(\mathbf{B}_{Q})_{b,n,r,d}\cdot(\bm{b}^{K}_{\text{cache}})_{b,m,d}.$
    This step captures the similarity between the feature parts of the query factors and the cached key factors.
  2. 2.
    Combine with query factor $\mathbf{A}_{Q}$ ($S^{(2)}$): The result $S^{(1)}$ is then combined with the head-specific query factors $\mathbf{A}_{Q}$. For each head $h$, the contributions from all query ranks $r$ are summed up.
    $S^{(2)}_{b,n,m,h}=\sum_{r=1}^{R_{Q}}(\mathbf{A}_{Q})_{b,n,h,r}\cdot S^{(1)}_{b,n,m,r}.$
    This yields an intermediate score for each query, cached key, and head.
  3. 3.
    Incorporate cached key factor $\mathbf{a}^{K}_{\text{cache}}$ for full logits ($\mathcal{L}$): The head-specific factor of the cached keys, $\mathbf{a}^{K}_{\text{cache}}$, is multiplied with $S^{(2)}$ to produce the final unscaled attention logits.
    $\mathcal{L}_{b,h,n,m}=S^{(2)}_{b,n,m,h}\cdot(\mathbf{a}^{K}_{\text{cache}})_{b,m,h}.$
    This logit $\mathcal{L}_{b,h,n,m}$ represents $s_{Q}^{-1}\cdot(\mathbf{Q}_{t}\mathbf{K}_{m}^{\top})_{h}$, i.e., the dot product between the query $\mathbf{Q}_{t}$ and the $m$-th cached key $\mathbf{K}_{m}$ for head $h$, scaled by $R_{Q}$ (since $s_{Q}=1/R_{Q}$). The ‘einsum‘ operation also rearranges dimensions to $(B,H,N,M)$ for the subsequent softmax operation.
  4. 4.
    Apply Scaling and Softmax ($\bm{\alpha}$): The logits $\mathcal{L}$ are scaled by the product of $s_{Q}$ (e.g., $1/R_{Q}$), $s_{K}$ (e.g., $1/R_{K}$, which is 1 if $R_{K}=1$), and $s_{\text{total}}$ (e.g., $1/\sqrt{D}$, where $D$ is the dimension used for dot products, typically $d_{h}$). Softmax is then applied across the $M$ cached key positions for each head independently.
    $\alpha_{b,h,n,m}=\operatorname{Softmax}_{m}\left(\mathcal{L}_{b,h,n,m}\cdot s_{Q}\cdot s_{K}\cdot s_{total}\right).$
    These are the attention probabilities.
  5. 5.
    Compute weighted sum with cached value factor $\mathbf{a}^{V}_{\text{cache}}$ ($\mathbf{O}^{(A)}$): The attention probabilities $\bm{\alpha}$ are used to weight the head-specific factors of the cached values, $\mathbf{a}^{V}_{\text{cache}}$.
    $\mathbf{O}^{(A)}_{b,n,m,h}=\alpha_{b,h,n,m}\cdot(\mathbf{a}^{V}_{\text{cache}})_{b,m,h}.$
    This step prepares the value components for the final aggregation. The ‘einsum‘ operation also reorders dimensions to $(B,N,M,H)$.
  6. 6.
    Incorporate cached value factor $\bm{b}^{V}_{\text{cache}}$ and apply final scaling ($\mathbf{O}$): Finally, the intermediate weighted value factors $\mathbf{O}^{(A)}$ are combined with the feature-specific factors of the cached values, $\bm{b}^{V}_{\text{cache}}$, by summing over the $M$ cached positions. The result is then scaled by $s_{V}$ (e.g., $1/R_{V}$, which is 1 if $R_{V}=1$).
    $\mathbf{O}_{b,n,h,e}=\left(\sum_{m=1}^{M}\mathbf{O}^{(A)}_{b,n,m,h}\cdot(\bm{b}^{V}_{\text{cache}})_{b,m,e}\right)\cdot s_{V}.$
[…]

### v5

@@ A.1 Direct computation in factor space
4:scales $s_{Q}{=}1/R_{Q}$, $s_{K}{=}1/R_{K}$, $s_{V}{=}1/R_{V}$; mask $\mathsf{mask}\in\{0,-\infty\}^{B\times T_{q}\times T_{k}}$
[…]
11:return $\mathrm{transpose}(\mathbf{O},~\text{"bhqe"}\to\text{"bqhe"})$
A.3 Complexity of the specialized path
Combining the terms above gives complexity $\mathcal{F}_{\text{TPA-spec}}=\Theta(BT^{2}R_{Q}R_{K}D)+\Theta(BHT^{2}R_{Q}R_{K})+\Theta(BHT^{2}R_{V}E)$, with the speed condition Eq. (A.4).
For a single query ($T_{q}{=}1$) against a cache of length $M$, the specialized FLOPs are
$\Theta(B\,M\,R_{Q}R_{K}\,D)\;+\;\Theta(B\,H\,M\,R_{Q}R_{K})\;+\;\Theta(B\,H\,M\,R_{V}\,E),$
while MHA uses $2\,\Theta(BHMD)$. This matches the asymptotics embodied in FlashTPA (Section 5) and explains the regimes where $R_{Q}{\ll}D$ and $R_{K}{=}R_{V}\in\{1,2\}$ yield the largest gains.
We apply the causal mask before softmax and use an online log-sum-exp update for numerical stability (as in FlashAttention). The intermediate $P\in\mathbb{R}^{B\times T_{q}\times T_{k}\times R_{Q}\times R_{K}}$ is evaluated blockwise in $T_{k}$ to keep peak memory linear in the block size; the same blocking naturally fuses with the masked softmax and the value aggregation step.
The constants $s_{Q},s_{K},s_{V}$ can be absorbed into either $\mathbf{A}_{(\cdot)}$ or $\mathbf{B}_{(\cdot)}$ at training time. We expose them explicitly only to make Eq. (A.4) transparent; The choice has no effect on softmax invariance or gradients.
The Triton kernel in Section 5 implements the blocked computation of $P$, the masked online softmax over $k$, and the fused value aggregation, mirroring Algorithm˜1. This avoids creating any $\mathbf{Q},\mathbf{K},\mathbf{V}$ or full $T_{q}{\times}T_{k}$ temporaries beyond working tiles.
Compared with $2\,\Theta(BHT^{2}D)$ for MHA, the specialized path improves with small $(R_{Q},R_{K},R_{V})$ and benefits further from pre-rotating $\mathbf{B}_{K}$ for RoPE (cf. Section 3.2), which removes per-step rotations in decoding. Practical speed also depends on tiling, memory bandwidth, and kernel fusion; our measured gains in Section 6.2 align with the regime predicted by Eq. (A.4).
A.4 Inference-time decoding cost across mechanisms
In autoregressive decoding, we generate the output for the current token $\mathbf{x}_{T}$ given cached keys and values from T1T{-}1 previous tokens. We analyze the FLOPs for computing the attention output for this single query token and use $M$ for the current cache length. For all mechanisms, we analyze the total Floating Point Operations (FLOPs) and the number of parameters in the attention layer, including the cost of projecting the current token’s hidden state $\mathbf{x}_{T}$ into its respective Query, Key, and Value representations. The parameter count formulas are taken from Table 1.
For Multi-Head Attention (MHA), with $H$ query heads and $H$ distinct Key/Value heads, the complexity is determined by the dot-product attention and value aggregation steps.
  • Projection: Projecting $\mathbf{x}_{T}$ to get a query, key, and value vector for each of the $H$ heads costs $\Theta(d_{\text{model}}Hd_{h})$.
  • Attention: The query vectors interact with a K/V cache of size $T$, costing $\Theta(MHd_{h})$.
  • Total MHA: The complexity is $\Theta(d_{\text{model}}Hd_{h}+MHd_{h})$.
Multi-Query Attention (MQA) uses $H$ query heads but shares a single Key/Value head ($H_{kv}=1$). The arithmetic complexity remains the same as MHA for the same number of query heads.
  • Projection: Projecting for $H$ query heads and 1 shared K/V head costs $\Theta(d_{\text{model}}(Hd_{h}+2d_{h}))$.
  • Attention: The interaction with the cache costs $\Theta(2MHd_{h})$.
  • Total MQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2)+MHd_{h})$.
Grouped-Query Attention (GQA) uses $H$ query heads and $N_{g}$ Key/Value head groups ($H_{kv}=N_{g}$). The arithmetic complexity is also identical to MHA.
  • Projection: Projecting for $H$ query heads and $N_{g}$ K/V head groups costs $\Theta(d_{\text{model}}(Hd_{h}+2N_{g}d_{h}))$.
  • Attention: The interaction with the cache costs $\Theta(2MHd_{h})$.
  • Total GQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2N_{g})+MHd_{h})$.
MQA and GQA significantly reduce the KV cache size and memory bandwidth compared to MHA. While the arithmetic FLOP count for the core attention computation (dot products and weighted sums) is $2MHd_{h}$ for all three (for fixed $H,d_{h}$), practical speedups for MQA/GQA arise from improved memory locality due to smaller K/V caches.
Multi-Head Latent Attention (MLA), as described in Appendix F.3, uses $H$ heads. For each head, the key $\mathbf{K}_{i}$ has dimension $d_{k}^{\prime}=d_{c}+d_{h}^{R}$, and the value $\mathbf{V}_{i}^{C}$ has dimension $d_{v}^{\prime}=d_{c}$.
  • Projection: Projecting $\mathbf{x}_{M}$ to get compressed and RoPE components for Q, K, and V costs $\Theta(d_{\text{model}}((d_{c}+d_{h}^{R})\cdot h+d_{c}+d_{h}^{R}))$.
  • Attention: The query vectors interact with a K/V cache of size $M$, costing $\Theta(MH(2d_{c}+d_{h}^{R}))$.
  • Total MLA: The complexity is $\Theta(d_{\text{model}}(d^{\prime}_{c}+2d_{c}+2d_{h}^{R})+MH(2d_{c}+d_{h}^{R}))$.
TPA. We use the FlashTPA Decoding algorithm (Algorithm 2) for FLOPs analysis, with $N=1$ query token, $T$ cached items, $D$ as feature dimension for $\mathbf{B}_{Q}/\bm{b}^{K}$ (typically $d_{h}$), and $E$ for $\bm{b}^{V}$ (typically $d_{h}$). For ranks $(R_{Q},R_{K},R_{V})$:
  • Projection: Projecting $\mathbf{x}_{M}$ to all Q, K, V factors costs $\Theta(d_{\text{model}}((R_{Q}+R_{K}+R_{V})(H+d_{h})))$.
  • Attention: The decoding algorithm’s interaction with the cache costs $\Theta(M[R_{K}(R_{Q}D+HR_{Q}+H)+R_{V}H(1+E)])$.
  • Total for TPA decoding: $\Theta(d_{\text{model}}(R_{Q}+R_{K}+R_{V})(H+d_{h})+M[R_{K}(R_{Q}D+HR_{Q}+H)+R_{V}H(1+E)])$.
Example Comparison I.
We compare the total Floating Point Operations (FLOPs) required to process a single token during autoregressive inference. This analysis separates the initial, constant projection cost from the attention cost, which scales linearly with the cache length $M$.
The following parameters are used for the comparison:
  • Model Dimension: $d_{\text{model}}=2048$
  • Heads: $H=32$
  • Head Dimension: $d_{h}=64$ (so $D=E=d_{h}$)
  • GQA Groups: $N_{g}=4$
  • MLA Dimensions: $d_{c}=256$, $d_{h}^{R}=32$, and $d^{\prime}_{c}=768$
MHA (16.8M parameters):
$\text{Parameters}=4d_{\text{model}}Hd_{h}=4\cdot 2048\cdot(32\cdot 64)\approx 16.8\times 10^{6}$
$\text{Projection}=3\cdot d_{\text{model}}\cdot H\cdot d_{h}=3\cdot 2048\cdot 32\cdot 64\approx 12.6\times 10^{6}$
$\text{Attention}=2\cdot M\cdot H\cdot d_{h}=4096M$
GQA ($N_{g}=4$, 9.4M parameters):
$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=2048\cdot 64\cdot(2\cdot 32+2\cdot 4)\approx 9.4\times 10^{6}$
$\text{Projection}=d_{\text{model}}(H+2N_{g})d_{h}=2048\cdot(32+8)\cdot 64\approx 5.2\times 10^{6}$
$\text{Attention}=2\cdot M\cdot H\cdot d_{h}=4096M$
MLA (9.8M parameters):
$\text{Parameters}=768(2048+2048+1024)+2048(32+2048)+256(2048+4096)\approx 9.8\times 10^{6}$
$\text{Projection}=d_{\text{model}}((d_{c}+d_{h}^{R})\cdot h+d_{c}+d_{h}^{R})=2048\cdot((256+32)\cdot 32+256+32)\approx 19.5\times 10^{6}$
$\text{Attention}=M\cdot H\cdot(2d_{c}+d_{h}^{R})=M\cdot 32\cdot(512+32)=17408M$
TPA ($R_{Q}=16,R_{K}=1,R_{V}=1$, 7.7M parameters):
$\text{Parameters}=d_{\text{model}}(16+1+1)(H+d_{h})+d_{\text{model}}Hd_{h}=2048(18)(96)+2048^{2}\approx 7.7\times 10^{6}$
$\text{Projection}=d_{\text{model}}(16+1+1)(H+d_{h})=2048\cdot(18)\cdot(96)\approx 3.5\times 10^{6}$
$\text{Attention}=M\cdot[1(1568)+1(2080)]=3648M$
TPA ($R_{Q}=16,R_{K}=2,R_{V}=2$, 8.1M parameters):
$\text{Parameters}=d_{\text{model}}(16+2+2)(H+d_{h})+d_{\text{model}}Hd_{h}=2048(20)(96)+2048^{2}\approx 8.1\times 10^{6}$
$\text{Projection}=d_{\text{model}}(16+2+2)(H+d_{h})=2048\cdot(20)\cdot(96)\approx 3.9\times 10^{6}$
$\text{Attention}=M\cdot[2(1568)+2(2080)]=7296M$
TPA ($R_{Q}=8,R_{K}=1,R_{V}=1$, 6.2M parameters):
$\text{Parameters}=d_{\text{model}}(8+1+1)(H+d_{h})+d_{\text{model}}Hd_{h}=2048(10)(96)+2048^{2}\approx 6.2\times 10^{6}$
$\text{Projection}=d_{\text{model}}(8+1+1)(H+d_{h})=2048\cdot(10)\cdot(96)\approx 2.0\times 10^{6}$
$\text{Attention}=M\cdot[1(800)+1(2080)]=2880M$
TPA ($R_{Q}=8,R_{K}=2,R_{V}=2$, 6.6M parameters):
$\text{Parameters}=d_{\text{model}}(8+2+2)(H+d_{h})+d_{\text{model}}Hd_{h}=2048(12)(96)+2048^{2}\approx 6.6\times 10^{6}$
$\text{Projection}=d_{\text{model}}(8+2+2)(H+d_{h})=2048\cdot(12)\cdot(96)\approx 2.4\times 10^{6}$
$\text{Attention}=M\cdot[2(800)+2(2080)]=5760M$
The analysis shows that TPA with low ranks offers a favorable trade-off. Reducing the query rank ($R_{Q}$) from 16 to 8 further decreases both the projection and attention costs, making the TPA ($R_{Q}{=}8,R_{K}{=}1,R_{V}{=}1$) configuration the most computationally efficient in this comparison. Increasing key/value ranks (e.g., to $R_{K}{=}2,R_{V}{=}2$) raises the attention cost linearly, remaining competitive with MHA for sufficiently long contexts where kernel fusion and blocking amortize memory traffic.
Example Comparison II.
We now repeat the analysis for a larger model configuration to observe how these trade-offs scale.
The following parameters for a larger model are used for this comparison:
  • Model Dimension: $d_{\text{model}}=4096$
  • Heads: $H=32$
  • Head Dimension: $d_{h}=128$ (so $D=E=d_{h}$)
  • GQA Groups: $N_{g}=4$
  • MLA Dimensions: $d_{c}=512$, $d_{h}^{R}=64$, and $d^{\prime}_{c}=1536$
MHA (67.1M parameters):
$\text{Parameters}=4d_{\text{model}}Hd_{h}=4\cdot 4096^{2}\approx 67.1\times 10^{6}$
$\text{Projection}=3\cdot 4096\cdot 32\cdot 128\approx 50.3\times 10^{6}$
$\text{Attention}=2\cdot M\cdot 32\cdot 128=8192M$
GQA ($N_{g}=4$, 37.7M parameters):
$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=4096\cdot 128\cdot(2\cdot 32+2\cdot 4)\approx 37.7\times 10^{6}$
$\text{Projection}=4096\cdot(32+8)\cdot 128\approx 21.0\times 10^{6}$
$\text{Attention}=2\cdot M\cdot 32\cdot 128=8192M$
MLA (39.1M parameters):
$\text{Parameters}=1536(4096+4096+2048)+4096(64+4096)+512(4096+8192)\approx 39.1\times 10^{6}$
$\text{Projection}=4096\cdot((512+64)\cdot 32+512+64)\approx 77.9\times 10^{6}$
$\text{Attention}=M\cdot 32\cdot(1024+64)=34816M$
TPA ($R_{Q}=16,R_{K}=1,R_{V}=1$, 28.6M parameters):
$\text{Parameters}=4096(18)(160)+4096^{2}\approx 28.6\times 10^{6}$
$\text{Projection}=4096\cdot(16+1+1)\cdot(32+128)\approx 11.8\times 10^{6}$
$\text{Attention}=M\cdot[1(2592)+1(4128)]=6720M$
TPA ($R_{Q}=16,R_{K}=2,R_{V}=2$, 29.9M parameters):
$\text{Parameters}=4096(20)(160)+4096^{2}\approx 29.9\times 10^{6}$
$\text{Projection}=4096\cdot(16+2+2)\cdot(32+128)\approx 13.1\times 10^{6}$
$\text{Attention}=M\cdot[2(2592)+2(4128)]=13440M$
TPA ($R_{Q}=8,R_{K}=1,R_{V}=1$, 23.3M parameters):
$\text{Parameters}=4096(10)(160)+4096^{2}\approx 23.3\times 10^{6}$
$\text{Projection}=4096\cdot(8+1+1)\cdot(32+128)\approx 6.6\times 10^{6}$
$\text{Attention}=M\cdot[1(1312)+1(4128)]=5440M$
TPA ($R_{Q}=8,R_{K}=2,R_{V}=2$, 24.6M parameters):
$\text{Parameters}=4096(12)(160)+4096^{2}\approx 24.6\times 10^{6}$
$\text{Projection}=4096\cdot(8+2+2)\cdot(32+128)\approx 7.9\times 10^{6}$
$\text{Attention}=M\cdot[2(1312)+2(4128)]=10880M$
For this larger configuration, TPA ($R_{Q}{=}8,R_{K}{=}1,R_{V}{=}1$) remains the clear leader in computational efficiency, with the lowest projection and attention costs. This highlights the value of tuning TPA ranks to balance expressiveness against compute.
Example Comparison III.
[…]
The following parameters are used for this comparison:
  • Model Dimension: $d_{\text{model}}=7168$
  • Heads: $H=64$
  • Head Dimension: $d_{h}=128$ (so $D=E=d_{h}$)
  • GQA Groups: $N_{g}=8$
  • MLA Dimensions: $d_{c}=512$, $d_{h}^{R}=64$, and $d^{\prime}_{c}=1536$
MHA (235M parameters):
$\text{Parameters}=4d_{\text{model}}Hd_{h}=4\cdot 7168\cdot 8192\approx 235\times 10^{6}$
$\text{Projection}=3\cdot 7168\cdot 64\cdot 128\approx 176.2\times 10^{6}$
$\text{Attention}=2\cdot M\cdot 64\cdot 128=16384M$
GQA ($N_{g}=8$, 132M parameters):
$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=7168\cdot 128\cdot(2\cdot 64+2\cdot 8)\approx 132\times 10^{6}$
$\text{Projection}=7168\cdot(64+16)\cdot 128\approx 73.4\times 10^{6}$
$\text{Attention}=2\cdot M\cdot 64\cdot 128=16384M$
MLA (101M parameters):
$\text{Parameters}=1536(7168+8192+4096)+7168(64+8192)+512(7168+16384)\approx 101\times 10^{6}$
$\text{Projection}=7168\cdot((512+64)\cdot 64+512+64)=268.4\times 10^{6}$
$\text{Attention}=M\cdot 64\cdot(2\cdot 512+64)=69632M$
TPA ($R_{Q}=16,R_{K}=1,R_{V}=1$, 83M parameters):
$\text{Parameters}=7168(18)(192)+7168\cdot 8192\approx 83.5\times 10^{6}$
$\text{Projection}=7168\cdot(16+1+1)\cdot(64+128)\approx 24.8\times 10^{6}$
$\text{Attention}=M\cdot[1(3136)+1(8256)]=11392M$
TPA ($R_{Q}=16,R_{K}=2,R_{V}=2$, 86.2M parameters):
$\text{Parameters}=7168(20)(192)+7168\cdot 8192\approx 86.2\times 10^{6}$
$\text{Projection}=7168\cdot(16+2+2)\cdot(64+128)\approx 27.5\times 10^{6}$
$\text{Attention}=M\cdot[2(3136)+2(8256)]=22784M$
TPA ($R_{Q}=8,R_{K}=1,R_{V}=1$, 72.5M parameters):
$\text{Parameters}=7168(10)(192)+7168\cdot 8192\approx 72.5\times 10^{6}$
$\text{Projection}=7168\cdot(8+1+1)\cdot(64+128)\approx 13.8\times 10^{6}$
$\text{Attention}=M\cdot[1(1600)+1(8256)]=9856M$
TPA ($R_{Q}=8,R_{K}=2,R_{V}=2$, 75.2M parameters):
$\text{Parameters}=7168(12)(192)+7168\cdot 8192\approx 75.2\times 10^{6}$
$\text{Projection}=7168\cdot(8+2+2)\cdot(64+128)\approx 16.5\times 10^{6}$
$\text{Attention}=M\cdot[2(1600)+2(8256)]=19712M$
At this very large scale, the cost of MHA projections becomes prohibitive. While MLA’s projection cost can be competitive, its attention cost scales with $(2d_{c}{+}d_{h}^{R})$ and exceeds MHA for long sequences. TPA with low ranks ($R_{Q}{=}8,R_{K}{=}1,R_{V}{=}1$) yields the lowest attention cost and a substantially smaller projection cost, strengthening its advantage as model size increases.
Appendix B More on FlashTPA Decoding Algorithm
In this section, we present FlashTPA for decoding in a hardware–friendly, numerically stable form and extend it to general key/value ranks $R_{K},R_{V}\geq 1$. The algorithm computes attention without materializing $\mathbf{Q},\mathbf{K},\mathbf{V}$ or the full $N{\times}M$ attention matrix, by (i) forming head‑shared feature–space dot products, (ii) mixing them with head‑specific factors to obtain logits as in Eq. (A.2), and (iii) aggregating values as in Eq. (A.3) in a single online softmax pass.
Notation and shapes.
We allow $N$ query positions but decoding uses N=1N{=}1. Let $B$ be batch, $M$ the cache length, $H$ heads, $R_{Q},R_{K},R_{V}$ ranks, and $D,E$ feature sizes (typically $D{=}E{=}d_{h}$). Inputs:
$\mathbf{A}_{Q}\!\in\!\mathbb{R}^{B\times N\times H\times R_{Q}},\;\mathbf{B}_{Q}\!\in\!\mathbb{R}^{B\times N\times R_{Q}\times D},\;\mathbf{A}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times H\times R_{K}},\;\mathbf{B}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{K}\times D},$
$\mathbf{A}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times H\times R_{V}},\;\mathbf{B}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{V}\times E}.$
We use scalings $s_{Q}{=}1/R_{Q}$, $s_{K}{=}1/R_{K}$, $s_{V}{=}1/R_{V}$, and $s_{\text{total}}{=}1/\sqrt{D}$. Let $\mathsf{mask}\!\in\!\{0,-\infty\}^{B\times N\times M}$ encode causality/padding. If RoPE pre‑rotation is used (Section 3.2), $\mathbf{B}_{K}^{\text{cache}}$ already includes positional phases; otherwise apply $\operatorname{RoPE}$ to $\mathbf{B}_{K}$ on load (no other change).
Algorithm 2 FlashTPA Decoding (general $R_{K},R_{V}$, masked, online‑LSE)
[…]
3:Initialize $\mathbf{y}\leftarrow 0^{B\times H\times N\times E}$, $\mathbf{lse}\leftarrow 0^{B\times H\times N}$, $\mathbf{m}\leftarrow(-\infty)^{B\times H\times N}$
4:for each cache block $m{:}m{+}\Delta m\!\leq\!M$ do
5:  Load $\mathbf{B}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{K}\times D}$, $\mathbf{A}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times H\times R_{K}}$
6:  Load $\mathbf{A}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times H\times R_{V}}$, $\mathbf{B}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{V}\times E}$, $\mathsf{mask}_{\text{blk}}\!\in\!\mathbb{R}^{B\times N\times\Delta m}$
7:  (1) Head‑shared feature dots:   $\mathbf{P}\leftarrow\mathrm{einsum}(\text{``bnrd,bmsd$\to$bnmrs''},~\mathbf{B}_{Q},~\mathbf{B}_{K,\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times N\times\Delta m\times R_{Q}\times R_{K}}$
8:  (2) Per‑head rank mixing to logits:
9:  $\mathcal{L}_{\text{blk}}\leftarrow(s_{\text{total}}s_{Q}s_{K})\cdot\mathrm{einsum}(\text{``bnhr,bmhs,bnmrs$\to$bhnm''},~\mathbf{A}_{Q},~\mathbf{A}_{K,\text{blk}},~\mathbf{P})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times\Delta m}$
10:  $\mathcal{L}_{\text{blk}}\leftarrow\mathcal{L}_{\text{blk}}+\mathrm{broadcast}(\mathsf{mask}_{\text{blk}})$
11:  (3) Online softmax update (no $\bm{\alpha}$ materialization):
12:  $\mathbf{m}_{\text{blk}}\leftarrow\max_{m}(\mathcal{L}_{\text{blk}})$; $\mathbf{p}_{\text{blk}}\leftarrow\exp(\mathcal{L}_{\text{blk}}-\mathbf{m}_{\text{blk}})$; $\mathbf{s}_{\text{blk}}\leftarrow\sum_{m}\mathbf{p}_{\text{blk}}$
13:  (4) Block value mixing:   $\mathbf{V}_{\text{blk}}\leftarrow\mathrm{einsum}(\text{``bmhu,bmue$\to$bhme''},~\mathbf{A}_{V,\text{blk}},~\mathbf{B}_{V,\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times H\times\Delta m\times E}$
14:  $\mathbf{y}_{\text{blk}}\leftarrow\mathrm{einsum}(\text{``bhnm,bhme$\to$bhne''},~\mathbf{p}_{\text{blk}},~\mathbf{V}_{\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times E}$
15:  (5) Fuse blocks with log‑sum‑exp:
16:  $\mathbf{m}_{\text{new}}\leftarrow\max(\mathbf{m},\mathbf{m}_{\text{blk}})$;   $\mathbf{y}\leftarrow\exp(\mathbf{m}-\mathbf{m}_{\text{new}})[...,None]\odot\mathbf{y}~+~\exp(\mathbf{m}_{\text{blk}}-\mathbf{m}_{\text{new}})[...,None]\odot\mathbf{y}_{\text{blk}}$
17:  $\mathbf{lse}\leftarrow\exp(\mathbf{m}-\mathbf{m}_{\text{new}})\odot\mathbf{lse}~+~\exp(\mathbf{m}_{\text{blk}}-\mathbf{m}_{\text{new}})\odot\mathbf{s}_{\text{blk}}$;  $\mathbf{m}\leftarrow\mathbf{m}_{\text{new}}$
18:end for
19:return $\mathbf{O}\leftarrow s_{V}\cdot\frac{\mathbf{y}}{\mathbf{lse}[...,None]}$ permuted to $(B,N,H,E)$
Step (1)–(2) implements Eq. (A.2); step (4)–(5) implements Eq. (A.3) while fusing the masked softmax with value aggregation via online log‑sum‑exp (as in FlashAttention), thereby avoiding any $\bm{\alpha}$ materialization. When $R_{K}{=}R_{V}{=}1$ the contractions reduce to the simpler einsums in Figure 2.
Complexity and working set.
Per block of $\Delta m$ cache items, the dominant FLOPs are
$\Theta(B\,N\,\Delta m\,R_{Q}R_{K}D)\;+\;\Theta(B\,H\,N\,\Delta m\,R_{Q}R_{K})\;+\;\Theta(B\,H\,\Delta m\,R_{V}E),$
[…]